What concentration of NaC2H3O2 must be added to 0.1 F HC3H3O2 to give a hydronium concentration of 2 x 10 to the negative 6 gfw per liter?


1
Expert's answer
2021-03-26T05:37:26-0400

Molar Mass of NaC2H3O2 = 82.0343 g/mol

Molar Mass of HC3H3O2 = 72.067

= HC3H3O2 0.1 F = Number of mfw/V

= 0.1× 72.067ml = 7.2067 mfw

Formalities of NaC2H3O2 = 0.1

0.1× 82.0343 = 8.2 mfw

Hydronium concentration = 1 ×10 ^ -7

2×10^-6/1×10^-7 = 2×10^1 = 20

8.2x/7.2 = 20

=18

= 18 g/L